# centertableaux

### v2

@@ 3.5.2. General case
$=\left(\sum_{\begin{subarray}{c}(\iota,\iota^{\prime},\iota_{1},\iota_{1}^{\prime}\cdots,\iota_{r},\iota_{r}^{\prime})\in\mathrm{YD}_{l}\times\mathrm{YD}_{l^{\prime}}\times\mathrm{YD}_{l_{1}}\times\mathrm{YD}_{l_{1}}\times\cdots\times\mathrm{YD}_{l_{r}}\times\mathrm{YD}_{l_{r}}\\ \iota\mathop{\sqcup}\limits^{r}\iota^{\prime}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}\mathop{\sqcup}\limits^{r}\iota_{r}=\iota({\mathcal{O}})\end{subarray}}\sharp\left({\mathrm{Irr}}_{{\boldsymbol{\lambda}}}\left({\mathrm{U}}(p^{\prime},q^{\prime});{\mathcal{O}}(\iota^{\prime})\right)\right)\right)$
@@
$=\left(\sum_{\begin{subarray}{c}(\iota,\iota^{\prime},\iota_{1},\iota_{1}^{\prime}\cdots,\iota_{r},\iota_{r}^{\prime})\in\mathrm{YD}_{l}\times\mathrm{YD}_{l^{\prime}}\times\mathrm{YD}_{l_{1}}\times\mathrm{YD}_{l_{1}}\times\cdots\times\mathrm{YD}_{l_{r}}\times\mathrm{YD}_{l_{r}}\\ \iota\mathop{\sqcup}\limits^{r}\iota^{\prime}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}\mathop{\sqcup}\limits^{r}\iota_{r}=\iota({\mathcal{O}})\end{subarray}}\sharp({\mathrm{Irr}}_{{\boldsymbol{\lambda}}}({\mathrm{U}}(p^{\prime},q^{\prime});{\mathcal{O}}(\iota^{\prime})))\right)$

### v3

@@ 3.5.2. General case
$=\sum_{\begin{subarray}{c}(\iota,\iota^{\prime},\iota_{1},\iota_{1}^{\prime}\cdots,\iota_{r},\iota_{r}^{\prime})\in\mathrm{YD}_{l}\times\mathrm{YD}_{l^{\prime}}\times\mathrm{YD}_{l_{1}}\times\mathrm{YD}_{l_{1}}\times\cdots\times\mathrm{YD}_{l_{r}}\times\mathrm{YD}_{l_{r}}\\ \iota\mathop{\sqcup}\limits^{r}\iota^{\prime}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}\mathop{\sqcup}\limits^{r}\iota_{r}=\iota({\mathcal{O}})\end{subarray}}\sharp\left({\mathrm{Irr}}_{{\boldsymbol{\lambda}}}\left({\mathrm{U}}(p^{\prime},q^{\prime});{\mathcal{O}}(\iota^{\prime})\right)\right)$
@@
$=\sum_{\begin{subarray}{c}(\iota,\iota^{\prime},\iota_{1},\iota_{1}^{\prime}\cdots,\iota_{r},\iota_{r}^{\prime})\in\mathrm{YD}_{l}\times\mathrm{YD}_{l^{\prime}}\times\mathrm{YD}_{l_{1}}\times\mathrm{YD}_{l_{1}}\times\cdots\times\mathrm{YD}_{l_{r}}\times\mathrm{YD}_{l_{r}}\\ \iota\mathop{\sqcup}\limits^{r}\iota^{\prime}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\iota_{1}\mathop{\sqcup}\limits^{r}\cdots\mathop{\sqcup}\limits^{r}\iota_{r}\mathop{\sqcup}\limits^{r}\iota_{r}=\iota({\mathcal{O}})\end{subarray}}\sharp({\mathrm{Irr}}_{{\boldsymbol{\lambda}}}({\mathrm{U}}(p^{\prime},q^{\prime});{\mathcal{O}}(\iota^{\prime})))$